7, Shannon deposits $20,000 in the bank today at 8% compounded annually. Starting three years from now, she makes equal withdrawals of SX per year for five years and then withdraws the entire amount ten years from now. If she withdraws $30,000 at EOY10, what is X?
Answer #1

Present value of withdrawals=cash withdrawals*Present value of discounting factor(rate%,time period)

20000=X/1.08^3+X/1.08^4+X/1.08^5+X/1.08^6+X/1.08^7+30000/1.08^10

20000=X[1/1.08^3+1/1.08^4+........+1/1.08^7]+30000/1.08^10

20000=X*3.423105313+30000*0.463193488

Hence X=(20000-30000*0.463193488)/3.423105313

which is equal to

=$1783.23(Approx).